OpenStudy (anonymous):

amy paid 54.88 for for a pair of running shoes that were 20%off how much was the regular price

OpenStudy (anonymous):

r=100%-20%=80% P=54.88 P=b*r 54.88=b(80%) 54.88/0.8=b b=68.6